# Text-to-Remote-Sensing-Image Retrieval beyond RGB Sources

### v1

@@ 4.2 The CLOSP architecture
The CLOSP architecture is a contrastive learning network whose training schema is depicted in Figure 5. It employs three different encoders, one for text, one for Sentinel-1 (SAR) data, and one for Sentinel-2 (optical) data. The encoders are trained to project their output representations into a shared latent space, aligning text with satellite imagery associated with the same item. Due to differences in satellite configurations, aligned images for the two visual modalities (i.e., SAR and optical) are rarely available. For example, the mean revisit time for a given area is 12 days for Sentinel-1 and 10 days for Sentinel-2. Thus, finding paired images for both modalities of the same area with a close timestamp, e.g., within the same day, is often not possible.
@@
During a training iteration, the text encoder generates a textual embedding for each of the $N$ elements ($M$ SAR and $M$ MSI) in the current batch. Then, each image element (either SAR or MSI) is processed by the corresponding vision encoder to produce visual embeddings. Finally, following the CLIP pretraining strategy and employing symmetric cross-entropy loss (see Equations 2 and 3), the textual and visual embeddings corresponding to the same item are simultaneously aligned (represented as white cells in the matrix in Figure 5) while being separated from all other elements (depicted as grey cells in Figure 5). The total loss $L$ is presented in Equation 4.
@@ 4.3 The GeoCLOSP architecture
Unlike natural images, geospatial imagery is inherently associated with geographic coordinates. Ideally, image embeddings within the same region should exhibit similarities. To also consider the geographical position of the satellite images, we propose GeoCLOSP, an extension of the CLOSP architecture tailored to align geographical coordinates with the visual modality. Inspired by SatCLIP (Klemmer et al., 2024), we utilize a Sinusoidal Representation Network (SIREN) (Sitzmann et al., 2020) and spherical harmonic (SH) positional encoding (Rußwurm et al., 2024) as a location encoder to embed the geographic coordinates associated with each image-text pair. The training schema, shown in Figure 5, builds upon the standard CLOSP alignment of text and visual modalities while introducing the alignment of image and location embeddings (see Equations 5 and 6). This dual alignment ensures that images with similar textual content share similar representations, and that images geographically close to each other also exhibit similarity. The loss $L_{g}$ follows the previous definition with an additional component as shown in Equation 7.
@@
$L_{g}=\frac{L_{\text{img}}+L_{\text{txt}}+L_{\text{loc}}+L_{\text{iloc}}}{4}$ (7)
Where $\mathbf{l}_{k}$ represents the embeddings for the $k$-th location associated with the corresponding image.
@@ 5.1 Baselines
We compare the performance of CLOSP with that of CLIP (Ilharco et al., 2021; Radford et al., 2021), SkyCLIP (Wang et al., 2024), RemoteCLIP (Liu et al., 2024), and SenCLIP (Jain et al., 2024) on the CrisisLandMark corpus. To the best of our knowledge, these models are the latest and best-performing T2RSIR models. We also fine-tune them on the training set of our dataset (hereafter we denote them as ¡ModelName¿-T). This provides an overview of the contribution of other spectral and radiometric information compared to RGB only. We also trained two specialized models (Text-SAR and Text-MS) to highlight the benefits of the unified space: we refer to the evaluation of these models together as BiCLIP. Additionally, we provide a dummy baseline for each set of experiments. All the specific settings regarding the analyzed models’ composition and efficiency are reported in Table 5.
@@ 5.3.1 Evaluation Metrics
We evaluate the retrieval performance in terms of Recall (R), Precision (P), and normalized Discounted Cumulative Gain (nDCG). We experiment with common cutoff levels $K\in\{10,50,100,1000\}$ to balance retrieval efficiency and common user needs. nDGC provides the most comprehensive evaluation for graded relevance. For the evaluation of precision and recall, we use a standard relevance threshold of 5 (i.e., IoU of 0.5). The random baseline performance is based on analytical computation: for each query, $R@K=K/|D|$, $P@K=R_{q}/|D|$ and $nDCG@K=R_{m}\frac{\sum_{i=1}^{K}1/log_{2}(i+1)}{IDCG@K}$, where $R_{m}$ is the average relevance score and $R_{q}$ is the number of relevant items.
@@ 5.3.3 Results
The text-to-image retrieval performance of our proposed models against several baselines is detailed in Table 6. Note that due to the large number of relevant items for some queries, a perfect Recall@1000 score is not always reachable. However, the relative differences demonstrate, in any case, the model’s capabilities. To confirm the observations, we performed a Friedman test followed by a Conover post-hoc test for multiple tests for statistical significance (p ¡ 0.01) (Demšar, 2006; Conover, 1998).
@@
Our CLOSP family of models significantly (according to the statistical test) outperforms all baselines (including BiCLIP) across every evaluated metric. The top-performing model, GeoCLOSP, achieves an nDCG@1000 of 57.76%, representing a nearly 20-point absolute improvement over the best baseline, SkyCLIP-T (37.88%).
This demonstrates the significant advantage of our unified architecture. It is also important to note that while CLOSP-VL employs a much larger backbone than the other variants, it provides only marginal gains. CLOSP-RN and CLOSP-VS do not exhibit a statistically significant difference from each other in terms of recall and precision.
@@
To visualize the ranking quality at various cutoffs, Figure 6 shows the nDCG and Precision curves for our CLOSP-RN and GeoCLOSP (which shares the same backbone) against the strongest baseline and the dummy predictor. The plots clearly illustrate that our models maintain a consistently higher performance across all values of k.
@@ 5.4.1 Results
The retrieval performance of our unified models compared to specialized baselines for each modality is presented in Table 7. The findings highlight the significant advantages of our joint training approach, particularly for the more challenging Sentinel-1 data.
@@
Conversely, for Sentinel-2 retrieval, our models remain highly competitive with the specialized baseline. CLOSP-RN, for instance, shows only a marginal performance decrease in nDCG@1000 compared to BiCLIP (54.80% vs 55.72%). A paired t-test confirms that while the gains on Sentinel-1 are statistically significant (p ¡ 0.01), the small performance decrease on Sentinel-2 is also significant, though minor in magnitude. This result demonstrates that our unified training strategy provides great gains for the SAR modality at a negligible cost of a minor drop in optical performance, confirming the value of the joint training paradigm.
@@ 5.5 Zero-shot classification
In this section, we analyze the performance of the CLOSP models in zero-shot multilabel classification over the whole corpus as performed in Radford et al. (2021) and Wang et al. (2024). Rather than measuring the model’s generalizability on unseen object categories, here we evaluate the “zero-shot” transfer capability on an unseen dataset (corpus without the training set), by employing the model without any additional training to classify the content. While some zero-shot evaluations focus on holding out entire classes, this approach is less suitable for foundational LULC classification. The selected classes form a comprehensive and interdependent set describing the terrestrial surface; training a model without knowledge of a fundamental class like “water” would inhibit its ability to learn related concepts such as “flooded vegetation”. Therefore, evaluating the transfer capability to unseen imagery is a more practical and meaningful test in this domain.
@@ 5.5.1 Settings
The multilabel classification is performed in a zero-shot fashion, leveraging the embeddings of the model. First, textual embeddings for each of the 12 keywords are generated using the model’s text encoder. Next, we compute the cosine similarity between each image embedding and each of the 12 keyword embeddings. To select a decision threshold for each model in a fair, zero-shot manner that accounts for their different score distributions, we adopted a global thresholding strategy. For a given model, we first computed the similarity scores for all 12 classes across all images in the corpus. The mean of this entire distribution of scores was then used as the single, model-specific classification threshold. A class was predicted as positive if its similarity score exceeded this global threshold. While more complex methods exist, this simple, dynamic threshold was chosen to maintain a strict zero-shot setting without requiring a validation set. We compare baselines and all CLOSP proposed models with a dummy classifier that always predicts the two most frequent classes, i.e., “crops” and “trees”. Additionally, when evaluating the BiCLIP solution, since each specialized model makes predictions for its image type without item overlapping, we merge the predictions of all dataset elements before computing the final performances. Finally, given the strong class imbalance, we evaluate the performance using macro-averaged precision (P), recall (R), and F1-score (F) over the classes.
@@ 5.5.2 Results
The results of the zero-shot classification are presented in Table 8. Our proposed CLOSP models demonstrate a clear advantage over existing baselines. Specifically, CLOSP-VS and CLOSP-RN achieve the highest performance with macro F1-scores of 41.82% and 41.56%, respectively, outperforming the best baseline model, i.e., SkyCLIP-T with an F1-score of 32.81%.
@@
To determine if differences are significant, we performed a Bayesian signed-rank test (Benavoli et al., 2014), which is more suitable than frequentist tests for comparisons with a small number of classes (12 labels) (Benavoli et al., 2017). CLOSP-VS, the best-performing model in terms of F1-score, is better than all baselines with high probability (¿ 0.9). CLOSP-RN achieves the highest recall, surpassing all models (probability larger than 0.9, except for SkyCLIP 0.88), proving its top-tier performance. CLOSP-VL, which achieves the best precision, also outperforms all baselines with high probability (¿ 0.9). Comparisons among the CLOSP variants were less conclusive, underscoring their similarities.
@@ 5.6.1 Settings
To analyze the relationship, we first drew two disjoint sets of 10,000 images each via uniform random sampling from the entire corpus. We then created 10,000 pairs by matching the first image from set A with the first from set B, the second with the second, and so on. For each of these pairs, we computed two distance values: the geographic distance using the Haversine formula and the embedding distance using the cosine distance (computed as 1 -- cosine similarity) between the embeddings. We analyze Pearson and Spearman correlations to understand both linear and monotonic relationships.
@@ 5.6.2 Results
To directly assess the impact of the location encoder, we compare the results from GeoCLOSP against its location-unaware counterpart, CLOSP-RN. For the CLOSP-RN model, we observe a negligible linear correlation (Pearson’s $r\approx-0.06$, $p<0.01$) and a weak monotonic correlation (Spearman’s $r_{s}\approx 0.13$, $p<0.01$). In contrast, while the GeoCLOSP model also shows no linear relationship (Pearson’s $r\approx-0.04$, $p<0.01$), it achieves a stronger monotonic correlation (Spearman’s $r_{s}\approx 0.34$, $p<0.01$). The relationship between geographical and embedding distances of the sampled sets of points is illustrated in Figure 7.
@@ 5.7 Single classes performance
The per-class performance is detailed in Table 9, revealing a clear trade-off between general semantic performance and specialization on geographic and crisis-related events. While SkyCLIP-T serves as a strong baseline, our statistical analysis confirms that both CLOSP-RN and GeoCLOSP significantly outperform it across the majority of classes for both metrics. The statistical analysis confirms that most observed performance differences are significant ($p<0.01$). The primary exceptions where no significant difference was found are:
  • 1.
@@ 6.2 The Synergy of a Unified Optical-SAR Latent Space
One of the most significant results is the demonstrated benefit of jointly training on optical and SAR data within a unified embedding space. As shown in our modality-specific evaluation (Table 7), the proposed approach improved retrieval performance for Sentinel-1 (SAR) imagery, a data source notoriously difficult for semantic interpretation due to its different imaging physics compared to optical sensors. This suggests a powerful knowledge transfer mechanism: the model leverages the clear semantic signatures from Sentinel-2’s multispectral bands (e.g., identifying vegetation or water) to learn corresponding structural and textural patterns in the Sentinel-1 SAR data, without any direct alignment strategy (there is no loss between the two vision encoders’ embeddings).
@@ 7 Conclusions and future work
CRediT authorship contribution statement
Daniele Rege Cambrin: Conceptualization, Methodology, Software, Investigation, Data Curation, Visualization, Writing - Original Draft; Lorenzo Vaiani: Visualization, Software, Conceptualization, Writing - Original Draft; Giuseppe Gallipoli: Validation, Writing - Original Draft, Data Curation; Luca Cagliero: Supervision, Writing - Review & Editing; Paolo Garza: Supervision, Writing - Review & Editing;
@@
Data availability
Appendix A Corine LC mapping

### v2

@@ 3.2 The CLOSP architecture
The CLOSP architecture is a contrastive learning network whose training schema is depicted in Figure˜5. It employs three different encoders, one for text, one for Sentinel-1 (SAR) data, and one for Sentinel-2 (optical) data. The encoders are trained to project their output representations into a shared latent space, aligning text with satellite imagery associated with the same item. Due to differences in satellite configurations, aligned images for the two visual modalities (i.e., SAR and optical) are rarely available. For example, the mean revisit time for a given area is 12 days for Sentinel-1 and 10 days for Sentinel-2. Thus, finding paired images for both modalities of the same area with a close timestamp, e.g., within the same day, is often not possible.
@@
During a training iteration, the text encoder generates a textual embedding for each of the $N$ elements ($M$ SAR and $M$ MSI) in the current batch. Then, each image element (either SAR or MSI) is processed by the corresponding vision encoder to produce visual embeddings. Finally, following the CLIP pretraining strategy and employing symmetric cross-entropy loss (see Equations˜2 and 3), the textual and visual embeddings corresponding to the same item are simultaneously aligned (represented as white cells in the matrix in Figure˜5) while being separated from all other elements (depicted as grey cells in Figure˜5). The total loss $L$ is presented in Equation˜4.
@@ 3.3 The GeoCLOSP architecture
Unlike natural images, geospatial imagery is inherently associated with geographic coordinates. Ideally, image embeddings within the same region should exhibit similarities. To also consider the geographical position of the satellite images, we propose GeoCLOSP, an extension of the CLOSP architecture tailored to align geographical coordinates with the visual modality. Inspired by SatCLIP (Klemmer et al., 2024), we utilize a Sinusoidal Representation Network (SIREN) (Sitzmann et al., 2020) and spherical harmonic (SH) positional encoding (RuSSwurm et al., 2024) as a location encoder to embed the geographic coordinates associated with each image-text pair. The training schema, shown in Figure˜5, builds upon the standard CLOSP alignment of text and visual modalities while introducing the alignment of image and location embeddings (see Equations˜5 and 6). This dual alignment ensures that images with similar textual content share similar representations, and that images geographically close to each other also exhibit similarity. The loss $L_{g}$ follows the previous definition with two additional components as shown in Equation˜7.
@@
$L_{g}=0.5\frac{L_{\text{img}}+L_{\text{txt}}}{2}+0.5\frac{L_{\text{loc}}+L_{\text{iloc}}}{2}$ (7)
Where $\mathbf{l}_{k}$ represents the embeddings for the $k$-th location associated with the corresponding image. An analysis on the impact of the weighting between the two components is reported in Appendix˜A.
@@ 4.1 Baselines
We compare the performance of CLOSP with that of CLIP (Ilharco et al., 2021; Radford et al., 2021), SkyCLIP (Wang et al., 2024), RemoteCLIP (Liu et al., 2024), SenCLIP (Jain et al., 2024), and Llama3-MS-CLIP (Marimo et al., 2025) (named Llama3-CLIP for simplicity) on the CrisisLandMark corpus. To the best of our knowledge, these models are the latest and best-performing T2RSIR models. We also fine-tune them on the training set of our dataset (hereafter we denote them as <ModelName>-T). This provides an overview of the contribution of other spectral and radiometric information compared to RGB only. We also trained two specialized models (Text-SAR and Text-MS) to highlight the benefits of the unified space: we refer to the evaluation of these models together as BiCLIP. We also trained a CLOPS-RGB model using an RGB encoder for Sentinel-2 and a False Color encoder for Sentinel-1 to additionally highlight the limitations of the interactions between RGB and SAR. Additionally, we provide a dummy baseline for each set of experiments. We provide CLOSP with three different backbones: ResNet-50 (CLOSP-RN), ViT-Small (CLOSP-VS), and ViT-Large (CLOSP-VL). All the specific settings regarding the analyzed models’ composition and efficiency are reported in Table˜5.
@@ 4.3.1 Evaluation Metrics
We evaluate the retrieval performance in terms of Recall (R), Precision (P), and normalized Discounted Cumulative Gain (nDCG) Manning et al. (2008). We experiment with common cutoff levels $K\in\{10,50,100,1000\}$ to balance retrieval efficiency and common user needs. nDGC provides the most comprehensive evaluation for graded relevance. For the evaluation of precision and recall, we use a standard relevance threshold of 5 (i.e., IoU of 0.5). The random baseline performance is based on analytical computation: for each query, $R@K=K/|D|$, $P@K=R_{q}/|D|$ and $nDCG@K=R_{m}\frac{\sum_{i=1}^{K}1/log_{2}(i+1)}{IDCG@K}$, where $R_{m}$ is the average relevance score and $R_{q}$ is the number of relevant items.
@@ 4.3.3 Results
The text-to-image retrieval performance of our proposed models against several baselines is detailed in Table˜6. Note that due to the large number of relevant items for some queries, a perfect Recall@1000 score is not always reachable. However, the relative differences demonstrate, in any case, the model’s capabilities. To confirm the observations, we performed a Friedman test followed by a Conover post-hoc test for multiple tests for statistical significance (p < 0.01) (Demšar, 2006; Conover, 1998).
@@
Our CLOSP family of models significantly (according to the statistical test) outperforms all baselines (including BiCLIP and CLOSP-RGB) across every evaluated metric. The top-performing model, GeoCLOSP, achieves an nDCG@1000 of 57.76%, representing a nearly 20-point absolute improvement over the best baseline, SkyCLIP-T (37.88%).
This demonstrates the significant advantage of our unified architecture. It is also important to note that while CLOSP-VL employs a much larger backbone than the other variants, it provides only marginal gains. CLOSP-RN and CLOSP-VS do not exhibit a statistically significant difference from each other in terms of recall and precision. Some visual examples are shown in Appendix˜B.
@@
To visualize the ranking quality at various cutoffs, Figure˜6 shows the nDCG and Precision curves for our CLOSP-RN and GeoCLOSP (which shares the same backbone) against the strongest baseline and the dummy predictor. The plots clearly illustrate that our models maintain a consistently higher performance across all values of k.
@@ 4.4.1 Results
The retrieval performance of our unified models compared to specialized baselines for each modality is presented in Table˜7. The findings highlight the significant advantages of our joint training approach, particularly for the more challenging Sentinel-1 data. CLOSP-RGB does not provide comparable results for either Sentinel-2 or Sentinel-1, so we focus our comparison on the best baseline, i.e., BiCLIP.
@@
Conversely, for Sentinel-2 retrieval, our models remain highly competitive with the specialized baseline. CLOSP-RN, for instance, shows only a marginal performance decrease in nDCG@1000 compared to BiCLIP (54.80% vs 55.72%). A paired t-test confirms that while the gains on Sentinel-1 are statistically significant (p < 0.01), the small performance decrease on Sentinel-2 is also significant, though minor in magnitude. This result demonstrates that our unified training strategy provides great gains for the SAR modality at a negligible cost of a minor drop in optical performance, confirming the value of the joint training paradigm.
@@ 4.5 Zero-shot classification
In this section, we analyze the performance of the CLOSP models in zero-shot multilabel classification over the whole corpus. Similarly to Radford et al. (2021) and Wang et al. (2024), we evaluate the “zero-shot” transfer capability of the model on an unseen dataset (corpus without the training set), instead of testing the model’s generalizability on unseen object categories. This tests the model’s generalization to new locations and scenes without any fine-tuning to assess how well the learned semantic representations can be used for direct classification. While some zero-shot evaluations focus on holding out entire classes, this approach is less suitable for foundational LULC classification. The selected classes form a comprehensive and interdependent set describing the terrestrial surface; training a model without knowledge of a fundamental class like “water” would inhibit its ability to learn related concepts such as “flooded vegetation”. Therefore, evaluating the transfer capability to unseen imagery is a more practical and meaningful test in this domain.
@@ 4.5.1 Settings
The multilabel classification is performed in a zero-shot fashion, leveraging the embeddings of the model. Given the text encoder $E_{t}$, the visual encoder $E_{v}$, and the set of 12 class (e.g; water, trees, burned area) keywords $C=\{c_{1},c_{2},...,c_{12}\}$, the task can be formalized in the following way. For each class keyword $c_{j}\in C$, we generate the class vector $v_{j}=E_{t}(c_{j})$. For each image $i$ in the corpus $D$, we generate the visual embeddings $u_{i}=E_{v}(i)$. We create a similarity matrix $S$ of shape $|D|\times 12$ computing the cosine similarity between each $u_{i}$ and $v_{j}$. A class $c_{j}$ is predicted as present in an image $i$ if its similarity score exceeds a threshold $t$.
To select a decision threshold $t$ for each model in a fair, zero-shot manner that accounts for their different score distributions, we adopted a global thresholding strategy. After the creation of $S$ for a specific model, the mean of this entire distribution of scores was then used as the single, model-specific classification threshold. While more complex methods exist, this simple, dynamic threshold was chosen to maintain a strict zero-shot setting without requiring a validation set.
We compare baselines and all CLOSP proposed models with a dummy classifier that always predicts the two most frequent classes, i.e., “crops” and “trees”. Additionally, when evaluating the BiCLIP solution, since each specialized model makes predictions for its image type without item overlapping, we merge the predictions of all dataset elements before computing the final performances. Finally, given the strong class imbalance, we evaluate the performance using macro-averaged precision (P), recall (R), and F1-score (F) over the classes.
@@ 4.5.2 Results
The results of the zero-shot classification are presented in Table˜8. Our proposed CLOSP models demonstrate a clear advantage over existing baselines. Specifically, CLOSP-VS and CLOSP-RN achieve the highest performance with macro F1-scores of 41.82% and 41.56%, respectively, outperforming the best baseline model, i.e., SkyCLIP-T with an F1-score of 32.81%.
@@
To determine if differences are significant, we performed a Bayesian signed-rank test (Benavoli et al., 2014), which is more suitable than frequentist tests for comparisons with a small number of classes (12 labels) (Benavoli et al., 2017). CLOSP-VS, the best-performing model in terms of F1-score, is better than all baselines with high probability (> 0.9). CLOSP-RN achieves the highest recall, surpassing all models (probability larger than 0.9, except for SkyCLIP 0.88), proving its top-tier performance. CLOSP-VL, which achieves the best precision, also outperforms all baselines with high probability (> 0.9). Comparisons among the CLOSP variants were less conclusive, underscoring their similarities.
@@ 4.6.1 Settings
To analyze the relationship, we first drew two disjoint sets of 10,000 images each via uniform random sampling from the entire corpus. We then created 10,000 pairs by matching the first image from set A with the first from set B, the second with the second, and so on. For each of these pairs, we computed two distance values: the geographic distance using the Haversine formula and the embedding distance using the cosine distance (computed as 1 - cosine similarity) between the embeddings. We analyze Pearson and Spearman correlations to understand both linear and monotonic relationships.
@@ 4.6.2 Results
To directly assess the impact of the location encoder, we compare the results from GeoCLOSP against its location-unaware counterpart, CLOSP-RN. For the CLOSP-RN model, we observe a negligible linear correlation (Pearson’s $r\approx-0.06$, $p<0.01$) and a weak monotonic correlation (Spearman’s $r_{s}\approx 0.13$, $p<0.01$). In contrast, while the GeoCLOSP model also shows no linear relationship (Pearson’s $r\approx-0.04$, $p<0.01$), it achieves a stronger monotonic correlation (Spearman’s $r_{s}\approx 0.34$, $p<0.01$). The relationship between geographical and embedding distances of the sampled sets of points is illustrated in Figure˜7.
@@ 4.7 Single classes performance
The per-class performance is detailed in Table˜9, revealing a clear trade-off between general semantic performance and specialization on geographic and crisis-related events. While SkyCLIP-T serves as a strong baseline, our statistical analysis confirms that both CLOSP-RN and GeoCLOSP significantly outperform it across the majority of classes for both metrics. The statistical analysis confirms that most observed performance differences are significant ($p<0.01$). The primary exceptions where no significant difference was found are:
@@ 5.2 The Synergy of a Unified Optical-SAR Latent Space
One of the most significant results is the demonstrated benefit of jointly training on optical and SAR data within a unified embedding space. As shown in our modality-specific evaluation (Table˜7), the proposed approach improved retrieval performance for Sentinel-1 (SAR) imagery, a data source notoriously difficult for semantic interpretation due to its different imaging physics compared to optical sensors. This suggests a powerful knowledge transfer mechanism: the model leverages the clear semantic signatures from Sentinel-2’s multispectral bands (e.g., identifying vegetation or water) to learn corresponding structural and textural patterns in the Sentinel-1 SAR data, without any direct alignment strategy (there is no loss between the two vision encoders’ embeddings).
@@ 6 Conclusions and future work
Code availability
The source code is available for downloading under the Apache-2.0 license at the link: https://github.com/DarthReca/closp.
Data availability
@@
Appendix A Analysis of Loss Weighting for GeoCLOSP
To analyze the balance between the semantic ($L_{txt-img}$) and geographic ($L_{img-loc}$) loss components, for simplicity, we can rewrite the loss shown in Equation˜7 as $L_{g}=\alpha L_{txt-img}+(1-\alpha)L_{img-loc}$, where $\alpha=0.5$ in this specific case. We tested $\alpha$ values of 0.25 and 0.75, comparing them to the baseline CLOSP-RN, which corresponds to $\alpha=1$, and GeoCLOSP ($\alpha=0.5$). The case where $\alpha=0$ (equivalent to SatCLIP) is omitted from the performance comparison as it discards the text query entirely, making it unsuitable for text-to-image retrieval. The results, reported in Table˜10, show that an equal weighting of $\alpha=0.5$ achieves the highest nDCG@1000 score of 57.76. This result outperforms the semantics-only baseline ($\alpha=1$) and suggests that incorporating geographic context is beneficial. Performance degrades as $\alpha$ shifts heavily towards either extreme, indicating that an even balance provides the optimal trade-off between semantic relevance and location-based alignment for this task.
Table 10: Impact of the loss weighting parameter $\alpha$ on text-to-image retrieval performance, measured by nDCG@1000. The parameter $\alpha$ balances the semantic loss ($L_{txt-img}$) and the geographic loss ($L_{img-loc}$).
Configuration $\alpha$ value nDCG@1000
CLOSP-RN (Semantics-only) 1.00 56.23
GeoCLOSP (High Semantics) 0.75 55.18
GeoCLOSP (Balanced) 0.50 57.76
GeoCLOSP (High Geography) 0.25 52.82
SatCLIP (Geography-only) 0.00 N/A*
    It discards the text alignment, making the model unable to perform the text-to-image retrieval task.
Appendix B Visual Examples of Retrieval
We reported some visual examples of retrieved images for increasingly complex queries in Figures˜8, 9 and 10. It is possible that nothing will be retrieved since vector databases (i.e., ChromaDB) use approximated KNN to navigate a large dataset. Similar images correspond to the same area collected at different timestamps.
CLOSP-RN 10 5 10 3 10 5
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
BiCLIP 5 5 5 2 3 7
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SkyCLIP-T 3 3 3 - - -
Refer to caption Refer to caption Refer to caption
Figure 8: Top-3 images by modality retrieved by the models for the provided query. We reported the RGB for Sentinel-2 and the VV polarization for Sentinel-1. Each image is coupled with the corresponding real relevance. Images are missing due to the use of approximated KNN in vector databases. Query: Flooded vegetation. Shrub and scrub
CLOSP-RN 8 8 3 8 5 6
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
BiCLIP 4 8 5 4 6 8
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SkyCLIP-T 6 2 4 - - -
Refer to caption Refer to caption Refer to caption
Figure 9: Top-3 images by modality retrieved by the models for the provided query. We reported the RGB for Sentinel-2 and the VV polarization for Sentinel-1. Each image is coupled with the corresponding real relevance. Images are missing due to the use of approximated KNN in vector databases. Query: Bare. Built. Crops. Trees
CLOSP-RN 7 5 7 4 4 3
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
BiCLIP 4 5 6 3 3 3
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SkyCLIP-T - - - 1 2 1
Refer to caption Refer to caption Refer to caption
Figure 10: Top-3 images by modality retrieved by the models for the provided query. We reported the RGB for Sentinel-2 and the VV polarization for Sentinel-1. Each image is coupled with the corresponding real relevance. Images are missing due to the use of approximated KNN in vector databases. Query: Bare. Burned area. Crops. Shrub and scrub. Snow and ice. Trees
Appendix C Corine LC mapping